Write explanation to the right of each step you would explain the following algebraic problem to learner
2.3*=162
3*=81
3*=34
X=4

Fractionize and solve the unknown in x square + 3x + 2
1
Expert's answer
2020-09-14T18:00:38-0400

We should solve an equation

"2\cdot 3^x = 162." First, we should divide each hand of the equation by 2, it helps us to obtain only the exponential term on the left hand of the equation.


"3^x=81." Next , we should determine, in which degree 3 becomes equal to 81. We know that "3^1=3, 3^2=9, 3^3=27, 3^4=81."


"3^x=3^4." We know that the exponential function is monotonic, so this equation has a single root, "x=4."



We should factorize "x^2+3x+2" . That means, we should transform it into "(x-x_1)(x-x_2)." If we expand this, we'll get "x^2-xx_1-xx_2+x_1x_2 = x^2-x(x_1+x_2)+x_1x_2."

Therefore, "x_1x_2 = 2, \;\; x_1+x_2 = -3." We may find such values, "x_1=-1, x_2 = -2."

Therefore, the factorization takes form "(x-(-1))(x-(-2)) = (x+1)(x+2)."